# Error Correction Code Transformer: From Non-Unified to Unified

### v2

@@ 1.2 Contributions
    High Computational Efficiency: The unified attention module in our proposed Transformer-based decoding architecture demonstrates linear computational complexity by eliminating the need to project the input into higher dimensions. This results in a significant reduction in complexity from the traditional $O(N^{2})$ to a more efficient $O(N)$, where $N$ represents the sequence length. Additionally, by employing our sparse mask, we further achieve an $86\%$ reduction in computational complexity. This efficiency gain enables faster processing and lower power consumption, making the architecture suitable for real-time applications in next-generation wireless communication systems.
@@
Notations: Bold lowercase and uppercase letters denote vectors (e.g., $\mathbf{v}$) and matrices (e.g., $\mathbf{M}$), respectively, while scalars use regular letters (e.g., $x$). $\mathbb{F}_{2}$ represents the Galois field of order 2, and $\mathbb{R}$, $\mathbb{N}$ denote the real and natural numbers. The transpose is indicated by $[\cdot]^{T}$, and $\mathcal{N}(\mu,\sigma^{2})$ denotes a Gaussian distribution with mean $\mu$ and variance $\sigma^{2}$. The function $\text{bipolar}(x)=1-2x$ maps $x\in{0,1}$ to bipolar values ${1,-1}$, with its inverse $\text{bin}(x)=\frac{1-x}{2}$ mapping ${1,-1}$ back to ${0,1}$. The operator $\text{size}(\cdot)$ denotes the number of elements in a vector or matrix, and $\text{sign}(\cdot)$ represents the sign function, defined as $\text{sign}(x)=1$ if $x>0$, 00 if $x=0$, and $-1$ if $x<0$. For a linear block code with code length $n$ and information bit length $k$, we denote it as $(n,k)$.
@@ 2.1 Vanilla Transformers
where $\text{head}_{i}=\text{Attn}(\textbf{{QW}}_{i}^{Q},\textbf{{KW}}_{i}^{K},\textbf{{VW}}_{i}^{V})$, with $\textbf{{Q}},\textbf{{K}},\textbf{{V}}\in\mathbb{R}^{N\times d_{k}}$ set to $\textbf{{X}}_{e}$ for self-attention, using trainable weights $\textbf{{W}}_{i}^{Q},\textbf{{W}}_{i}^{K},\textbf{{W}}_{i}^{V}\in\mathbb{R}^{d_{k}\times d_{k}}$ and $\textbf{{W}}^{O}\in\mathbb{R}^{Hd_{k}\times Hd_{k}}$, as depicted in Figure 2.
@@
Finally, residual connections and layer normalization link the MHA and FFN layers:
@@ 2.2 Forward Error Correction Codes
[…]
@@
y~=[|y|,s($\textbf{{y}}\right)\right]$ = [ | y | , s ( y ) ] (7)
@@
$\hat{\textbf{{x}}}=\text{bin}\left(\text{sign}\left({\textbf{{y}}}\cdot{\hat{\textbf{{z}}}}\right)\right)$ (8)
@@ 3.1 Shared Unified Attention
The vanilla self-attention module can be viewed as establishing correlations among different input features within a data sample while neglecting the intrinsic relationships between different samples. This limitation hinders its ability to learn the distinctiveness of features between various code types in channel decoding tasks. Furthermore, research in [30] suggests that the input of the softmax function is probabilistically sparse, indicating that only a few weights dominate. This implies that an $N\times N$ self-attention matrix is not strictly necessary. This sparsity suggests that a low-rank approximation could effectively capture the dominant correlations without sacrificing performance.
@@
To validate this, we trained an ECCT model (with parameters $L=6$, $H=8$, $d_{k}=32$, $d_{f}=1024$) on a variety of linear block codes, including Polar(32, 16), Polar(64, 32), Polar(128, 86), LDPC(49, 24), LDPC(121, 60), LDPC(121, 80), BCH(31, 16), BCH(63, 36), and BCH(127, 120). We analyzed the rank distribution of the $\textbf{{Q}}\cdot\textbf{{K}}^{T}$ matrix across 1000 batches for the 6-layer Transformer encoder. As shown in Figure 5, the rank ranges from 11 to 32, consistently indicating a low-rank structure across all layers and code types.
Based on the aforementioned considerations, we have developed a novel, low-complexity Unified Attention Module (UAM) to tackle channel decoding tasks involving multiple code types, lengths, and rates. We introduced learnable memory components, $\textbf{{A}}_{l}$ (attention) and $\textbf{{V}}_{l}$ (values), which play the role of learning features from the entire training dataset. The mathematical expression is given by:
@@
where $\textbf{{X}}\in\mathbb{R}^{N\times d_{k}}$ is the input of attention layer, $\textbf{{A}}_{l},\textbf{{V}}_{l}\in\mathbb{R}^{N\times d_{l}}$ are trainable parameters, and $d_{l}$ is an adjustable hyperparameter controlling the low-rank approximation.
In order to gain a better understanding of the proposed UAM, we rewrite equation (2) as follows:
@@ 6 $\textbf{{A}}=\text{Softmax}(\textbf{{A}}_{l},\text{dim}=-1)$  # $\text{shape}=(B,1,N,d_{l})$
7 $_{o}=\textbf{{A}}\cdot\left(\textbf{{V}}_{l}^{T}\cdot\textbf{{X}}\right)$ X )  # $\text{shape}=(B,H,N,d_{k})$
@@ 10 Output: $\textbf{{X}}_{o}$  # $\text{shape}=(B,N,d_{h})$
By comparing equations (9) and (10), the computational complexity of the vanilla self-attention module is $O(N^{2}\times d_{k}+N\times d_{k}^{2})$, whereas the proposed UAM reduces it to $O(N\times d_{l}\times d_{k})$. Importantly, the omission of the Q and K projections in the vanilla Transformer leads to significant parameter savings. As demonstrated in subsequent experiments, a $d_{l}$ value comparable to $n-k$ is sufficient to achieve a favorable bit error rate. This is supported by the fact that the maximum rank of a parity-check matrix $\textbf{H}\in\{0,1\}^{(n-k)\times n}$ is $n-k$, providing a theoretical basis for the observed performance. It is also important to note that both ECCT and FECCT are based on the vanilla Transformer architecture, differing only in their self-attention mask designs.
The proposed multi-head unified-attention and corresponding Python-style pseudocode are illustrated in Figure 6 and Algorithm 1, respectively. As depicted in Figure 6, thanks to the sparsity of the attention scores, multiple heads are able to share the same attention $\textbf{{A}}_{l}$, resulting in a more cohesive decoding architecture and further reducing the complexity of the neural network. This approach contrasts with vanilla self-attention mechanisms, in which each head independently computes its scores using separate learned weights.
@@ 3.2 Sparse Masked Unified Attention
[…]
@@
Inspired by this, we explore a mask structure leveraging the parity-check constraints. Let $\textbf{{A}}_{l}$ be a matrix in $\mathbb{R}^{N\times d_{l}}$, where $N$ is defined as $2n-k$ and $d_{l}$ as $n-k$. We then construct the extended parity-check matrix $\overline{\textbf{H}}$ as follows:
@@ 4 $\overline{\textbf{H}}[n:2n-k,0:n-k]=\text{eye}(n-k)$
5 Output: ()(¬$(-\infty)\cdot(\neg\mathrm{~{}\overline{\textbf{H}}})$
This illustration demonstrates that the application of the mask results in sparser attention, a phenomenon particularly pronounced in low-density parity-check codes. Most importantly, the complexity of the UAM is now reduced to the density of the code $O(N\times d_{l}\times d_{k}\times H_{d})$. Our experimental results indicate that the use of this mask has successfully reduced the computational complexity of the attention mechanism by an average of 86%. Furthermore, we provide a summary of the Python-style sparse mask construction within Algorithm 2.
@@ 3.3 Architecture and Training Methodology
In Figure 8, the output y from the AWGN channel passes through the pre-processing module to yield $N=2n-k$ input elements for the neural network. The decoder is constructed by concatenating $L$ Transformer encoders, which consist of unified-attention and feed-forward layers, with normalization layers in between. The output module is characterized by two fully connected layers. The initial layer compresses the multi-head output into a one-dimensional vector of size $2n-k$, with the subsequent layer transforming it into an $n$-dimensional vector representing the softly decoded noise. Finally, the post-processing module acts on the estimated noise to derive the transmitted original codeword.
We employ binary cross-entropy as the loss function with the objective of learning to predict the multiplicative noise z. The corresponding target binary multiplicative noise is denoted as z~=bin(sign($\textbf{{z}}))$ = bin ( sign ( z ) ). Therefore, the loss calculation for a received individual codeword y is given by:
@@
[…]
@@
where $\textbf{0}_{c}$ and $\textbf{0}_{s}$ represent zero vectors with sizes $N_{\text{max}}-\text{size}(\textbf{{y}})$ and $S_{\text{max}}-\text{size}(s(\textbf{{y}}))$, respectively. Through this unit, the Transformer neural decoder can accommodate any codeword within the maximum length, allowing for mixed training in a unified manner.
@@ 4 Experiments
Table 1 presents the simulation results, specifically showing the negative natural logarithm of the BER at $E_{b}/N_{0}$ values of 4, 5, and 6 dB. It is noted that higher values in this context indicate better performance, with the best results for each SNR highlighted in bold. Additional BER plots for Polar and BCH codes are provided in Figure 9 and Figure 10, respectively. Simulation results demonstrate that our proposed UECCT outperforms both the individually trained ECCT [24] and the hybrid-trained FECCT [25] across the majority of codewords, validating the effectiveness of the proposed architecture. Notably, UECCT even surpasses FECCT’s performance with an embedding length of 128 when operating at a more compact embedding length of 64, further highlighting the efficiency and power of our design. The superior performance stems from its ability to leverage a shared attention mechanism across heads, as detailed in Section 3, combined with sparsity-aware masking that reduces computational complexity while preserving decoding accuracy.

### v3

@@ 1.2 Contributions
    High Computational Efficiency: The unified attention module in our proposed Transformer-based decoding architecture demonstrates linear computational complexity by eliminating the need to project the input into higher dimensions. This results in a significant reduction in complexity from the traditional $\mathcal{O}(N^{2})$ to a more efficient $\mathcal{O}(N)$, where $N$ represents the sequence length. Additionally, by employing our sparse mask, we further achieve an $86\%$ reduction in computational complexity. This efficiency gain enables faster processing and lower power consumption, making the architecture suitable for real-time applications in next-generation wireless communication systems.
@@
Notations: Bold lowercase and uppercase letters denote vectors (e.g., $\mathbf{v}$) and matrices (e.g., $\mathbf{M}$), respectively, while scalars use regular letters (e.g., $x$). $\mathbb{F}_{2}$ represents the Galois field of order 2, and $\mathbb{R}$, $\mathbb{N}$ denote the real and natural numbers. The transpose is indicated by $\cdot]^{T}$, and $\mathcal{N}(\mu,\sigma^{2})$ denotes a Gaussian distribution with mean $\mu$ and variance $\sigma^{2}$. The function $\text{bipolar}(x)=1-2x$ maps $x\in{0,1}$ to bipolar values 1,1{1,-1}, with its inverse $\text{bin}(x)=\frac{1-x}{2}$ mapping 1,1{1,-1} back to 0,1{0,1}. The operator $\text{size}(\cdot)$ denotes the number of elements in a vector or matrix, and $\text{sign}(\cdot)$ represents the sign function, defined as $\text{sign}(x)=1$ if $x>0$, 0 if $x=0$, and 1-1 if $x<0$. For a linear block code with code length $n$ and information bit length $k$, we denote it as $(n,k)$.
@@ 2.1 Vanilla Transformers
where $\text{head}_{i}=\text{Attn}(\textbf{{X}}_{Q}\textbf{{W}}_{i}^{Q},\textbf{{X}}_{K}\textbf{{W}}_{i}^{K},\textbf{{X}}_{V}\textbf{{W}}_{i}^{V})$, with $\textbf{{X}}_{Q},\textbf{{X}}_{K},\textbf{{X}}_{V}\in\mathbb{R}^{N\times d_{k}}$ set to $\textbf{{X}}_{e}$ for self-attention, using trainable weights $\textbf{{W}}_{i}^{Q},\textbf{{W}}_{i}^{K},\textbf{{W}}_{i}^{V}\in\mathbb{R}^{d_{k}\times d_{k}}$ and $\textbf{{W}}^{O}\in\mathbb{R}^{Hd_{k}\times Hd_{k}}$, as depicted in Figure 2.
@@
Finally, residual connections and layer normalization link the MHA and FFN layers to produce the output $\textbf{{X}}_{2}$ of a single-layer Transformer encoder:
@@ 2.2 Forward Error Correction Codes
The encoder maps a message vector $\textbf{{m}}\in\{0,1\}^{k}$ to a codeword $\textbf{{x}}\in\{0,1\}^{n}$ via $\textbf{{x}}=\textbf{{m}}\cdot\textbf{G}$, where $\textbf{H}\cdot\textbf{{x}}=\textbf{0}$. This codeword is modulated using Binary Phase Shift Keying (BPSK), defined as $\text{BPSK}(x)=1-2x$, yielding $\textbf{{x}}_{s}$, which is transmitted over Binary-Input Discrete Memoryless Channels (B-DMCs, e.g., AWGN), producing $\textbf{{y}}=\textbf{{x}}_{s}+\textbf{{n}}$, with $\textbf{{n}}\sim\mathcal{N}(0,\sigma^{2})$.
@@
$\widetilde{\textbf{{y}}}=\left[\lvert\textbf{{y}}\rvert,\text{s}\left(\textbf{{y}}\right)\right]$ (7)
where, $\left[\cdot,\cdot\right]$ denotes the concatenation of vectors, $\lvert\textbf{{y}}\rvert$ signifies the absolute value of y and $\text{s}(\textbf{{y}})=\textbf{H}\cdot\text{bin}(\text{sign}(\textbf{{y}}))\in\{0,1\}^{n-k}$ is the syndrome from hard-decoded y.
@@ 3.1 Shared Unified Attention
The vanilla self-attention mechanism in Transformer architectures establishes dense correlations among all input features within a sequence, resulting in a computational complexity of $\mathcal{O}(N^{2}\cdot d_{k}+N\cdot d_{k}^{2})$, where $N$ is the sequence length and $d_{k}$ is the embedding dimension. This approach, while effective for capturing long-range dependencies, is computationally intensive and overlooks the intrinsic relationships across different samples, particularly in channel decoding tasks involving diverse linear block codes. Moreover, prior studies [30] suggest that the input to the softmax function in self-attention is probabilistically sparse, with only a few weights dominating the attention distribution. This sparsity implies that a low-rank approximation could effectively capture the dominant correlations, reducing redundancy without compromising performance.
@@
To investigate this hypothesis, we trained an ECCT model [24] with parameters $L=6$ (number of Transformer encoder layers), $H=8$, $d_{k}=\text{max}(2n-k)=182$, and $d_{f}=2912$ on a diverse set of linear block codes, including Polar(32,16), Polar(64,32), Polar(128,86), LDPC(49,24), LDPC(121,60), LDPC(121,80), BCH(31,16), BCH(63,36), and BCH(127,120). We analyzed the rank distribution of the $\textbf{{Q}}\cdot\textbf{{K}}^{T}$ matrix across 1000 batches for the six-layer Transformer encoder. As illustrated in Figure 5, the rank ranges from 12 to 181, with over 90% of values below 69, consistently indicating a low-rank structure across all layers and code types. This observation suggests that the high-dimensional self-attention matrix can be approximated with a lower-rank representation, offering a pathway to reduce computational overhead.
Figure 6: Distribution of Jensen-Shannon divergence between attention head pairs across 6 layers, showing high similarity (JSD << 0.17).
Furthermore, the conventional multi-head self-attention mechanism computes independent attention scores for each head using distinct weight matrices $\textbf{{W}}_{i}^{Q}$, $\textbf{{W}}_{i}^{K}$, and $\textbf{{W}}_{i}^{V}$. However, we hypothesize that the attention distributions across heads may exhibit high similarity, particularly in channel decoding, where the sparse structure of the parity-check matrix $\mathbf{H}$ constrains bit interactions. To quantify this, we employed the Jensen-Shannon Divergence (JSD) [31], a symmetric measure of similarity between probability distributions, defined as:
$\text{JSD}(\mathbf{A}_{i}[k]\|\mathbf{A}_{j}[k])=\frac{1}{2}\left[D_{KL}(\mathbf{A}_{i}[k]\|\mathbf{M}[k])+D_{KL}(\mathbf{A}_{j}[k]\|\mathbf{M}[k])\right]$ (9)
where $k$ is the row index, $\mathbf{A}_{i}$ and $\mathbf{A}_{j}$ $\in\mathbb{R}^{N\times N}$are the softmax outputs of attention heads $i$ and $j$, $\mathbf{M}[k]=\frac{1}{2}(\mathbf{A}_{i}[k]+\mathbf{A}_{j}[k])$ is the average distribution, and $D_{KL}(\mathbf{P}\|\mathbf{Q})=\sum_{l=1}^{N}\mathbf{P}[l]\log\frac{\mathbf{P}[l]}{\mathbf{Q}[l]}$ is the Kullback-Leibler divergence [32]. Using the same simulation setup as the low-rank validation ($L=6$, $H=8$, $d_{k}=182$, $d_{f}=2912$), we computed the per-row JSD for all head pairs across all rows, averaging the results. As shown in Figure 6, JSD values are consistently below 0.17 across all layers, with a mean of 0.13, indicating strong alignment in attention distributions. This similarity, driven by the sparse structure of $\mathbf{H}$, supports the feasibility of sharing a common attention matrix across heads.
Motivated by these findings, we propose a novel Unified Attention Module (UAM) to address the computational inefficiency and lack of cross-code generalization in traditional self-attention. The UAM introduces learnable memory components, $\textbf{{A}}_{l}$ and $\textbf{{V}}_{l}\in\mathbb{R}^{N\times d_{l}}$, where $d_{l}$ is a hyperparameter controlling the low-rank approximation. The UAM is defined as:
@@
where $\textbf{{X}}\in\mathbb{R}^{N\times d_{k}}$ is the input to the attention layer. In contrast, the conventional self-attention is expressed as:
@@
By eliminating the projections $\textbf{{W}}^{\textit{Q}}$ and $\textbf{{W}}^{\textit{K}}$, the UAM reduces the computational complexity from $\mathcal{O}(N^{2}\cdot d_{k}+N\cdot d_{k}^{2})$ to $\mathcal{O}(N\cdot d_{l}\cdot d_{k})$, where $d_{l}\ll N$. The shared attention matrix $\mathbf{A}_{l}$ is applied across all heads, leveraging the observed head similarity to foster a cohesive decoding architecture. This design not only reduces the parameter count but also enables the model to learn shared features across diverse code types, facilitating a code-agnostic framework.
@@ 6 $\textbf{{A}}=\text{Softmax}(\textbf{{A}}_{l},\text{dim}=-1)$  # $\text{shape}=(B,1,N,d_{l})$
7 $\textbf{{X}}_{o}=\textbf{{A}}\cdot\left(\textbf{{V}}_{l}^{T}\cdot\textbf{{X}}\right)$  # $\text{shape}=(B,H,N,d_{k})$
@@ 10 Output: $\textbf{{X}}_{o}$  # $\text{shape}=(B,N,d_{h})$
The architecture of the proposed multi-head UAM is depicted in Figure 7, with the corresponding pseudocode provided in Algorithm 1. The shared attention mechanism, combined with the low-rank approximation, significantly reduces model complexity while preserving decoding accuracy, as validated in subsequent experiments. This approach marks a significant advancement over the vanilla Transformer-based ECCT [24], which lacks cross-code unification.
@@ 3.2 Sparse Masked Unified Attention
The design of the mask function $\text{M}(\textbf{H})$ aims to enable the neural network to capture the relationship between the received signal $\textbf{{y}}\in\mathbb{R}^{n}$ and its syndrome $\text{s}(\textbf{{y}})=\textbf{H}\cdot\overline{\textbf{{y}}}\in\{0,1\}^{n-k}$, where $\overline{\textbf{{y}}}=\text{bin}(\text{sign}(\textbf{{y}}))$ is the hard-decoded binary vector derived from y. For a valid codeword y (i.e., noise-free or correctly decoded), the orthogonality condition $\textbf{H}\cdot\overline{\textbf{{y}}}=\textbf{0}$ holds over $\mathbb{F}_{2}$, yielding $\text{s}(\textbf{{y}})=\textbf{0}$. To embed this relationship within the attention mechanism, we define an extended parity-check matrix $\overline{\textbf{H}}\in\mathbb{R}^{(2n-k)\times(n-k)}$ satisfying:
@@
Inspired by this, we explore a mask structure leveraging the parity-check constraints. Let $\textbf{{A}}_{l}$ be a matrix in $\mathbb{R}^{N\times d_{l}}$, where $N$ is defined as 2nk2n-k and $d_{l}$ as nkn-k. The theoretical foundation for setting $d_{l}\approx n-k$ stems from the maximum rank of the parity-check matrix $\textbf{H}\in\mathbb{F}_{2}^{(n-k)\times n}$, which is nkn-k. We then construct the extended parity-check matrix $\overline{\textbf{H}}$ as follows:
@@ 4 $\overline{\textbf{H}}[n:2n-k,0:n-k]=\text{eye}(n-k)$
5 Output: $(-\infty)\cdot(\neg\mathrm{~\overline{\textbf{H}}})$
This illustration demonstrates that the application of the mask results in sparser attention, a phenomenon particularly pronounced in low-density parity-check codes. Most importantly, the complexity of the UAM is now reduced to the density of the code $\mathcal{O}(N\times d_{l}\times d_{k}\times H_{d})$. Our experimental results indicate that the use of this mask has successfully reduced the computational complexity of the attention mechanism by an average of 86%. Furthermore, we provide a summary of the Python-style sparse mask construction within Algorithm 2. It should be noted that, in Application-Specific Integrated Circuit (ASIC) implementations, computations involving the masked parts can be directly skipped to enhance efficiency.
@@ 3.3 Architecture and Training Methodology
In Figure 9, the output y from the AWGN channel passes through the pre-processing module to yield N=2nkN=2n-k input elements for the neural network. The decoder is constructed by concatenating $L$ Transformer encoders, which consist of unified-attention and feed-forward layers, with normalization layers in between. The output module is characterized by two fully connected layers. The initial layer compresses the multi-head output into a one-dimensional vector of size 2nk2n-k, with the subsequent layer transforming it into an $n$-dimensional vector representing the softly decoded noise. Finally, the post-processing module acts on the estimated noise to derive the transmitted original codeword.
We employ binary cross-entropy as the loss function with the objective of learning to predict the multiplicative noise z. The corresponding target binary multiplicative noise is denoted as $\tilde{\textbf{{z}}}=\text{bin}(\text{sign}(\textbf{{z}}))$. Therefore, the loss calculation for a received individual codeword y is given by:
@@
To accommodate codewords of varying lengths within a unified architecture, we have designed a standardization unit that performs padding on $\lvert\textbf{{y}}\rvert$ and $s(\textbf{{y}})$ as described in Equation (7). Specifically, after receiving the channel output, we pad different codewords and their corresponding syndromes with zeros up to the maximum length. Assuming that the maximum length of the codeword is $N_{\text{max}}$ and the maximum length of the syndrome is $S_{\text{max}}$, the dataset is standardized according to:
@@
where $\textbf{0}_{c}$ and $\textbf{0}_{s}$ represent zero vectors with sizes $N_{\text{max}}-\text{size}(\textbf{{y}})$ and $S_{\text{max}}-\text{size}(s(\textbf{{y}}))$, respectively. Through this unit, the Transformer neural decoder can accommodate any codeword within the maximum length, allowing for mixed training in a unified manner. In ASIC implementations, computations for zero-padded parts can be skipped to improve efficiency, similar to the sparse-masked attention.
@@ 4 Experiments
Figure 11: BER comparison of ECCT and proposed UECCT for LDPC codes.
@@
Table 1 presents the simulation results, specifically showing the negative natural logarithm of the BER at $E_{b}/N_{0}$ values of 4, 5, and 6 dB. It is noted that higher values in this context indicate better performance, with the best results for each SNR highlighted in bold. Additional BER plots for Polar, LDPC and BCH codes are provided in Figure 10, 11 and Figure 12, respectively. Simulation results demonstrate that our proposed UECCT outperforms both the individually trained ECCT  [24] and the hybrid-trained FECCT  [25] across the majority of codewords, validating the effectiveness of the proposed architecture. Notably, UECCT even surpasses FECCT’s performance with an embedding length of 128 when operating at a more compact embedding length of 64, further highlighting the efficiency and power of our design. The superior performance stems from its ability to leverage a shared attention mechanism across heads, as detailed in Section 3, combined with sparsity-aware masking that reduces computational complexity while preserving decoding accuracy.